Recent Log Entries

0L03-4 Tuner Issues

Log Number: 3819022Logbook(s): Entry Makers: aikenBody: 

Detune angle started flopping back and forth at large values. Cavity is tuning back and forth, but it is not creating any stability in the detune angle. Energy is significantly low coming out of the INJ as well.

Dipole Comm Errors

Log Number: 3819019Logbook(s): Entry Makers: aikenBody: 

Still investigating. Was paired with a familiar MZA8R04 mismatch.

Operator SIGN IN

Log Number: 3819017Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Deir
DATE:             08/01/2020
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Downtime Incident CLOSED: 1L14-4 recovery

Log Number: 3819015Logbook(s): Body: IDDescriptionPeriodCause8267Title:
1L14-4 recovery

Summary:
EES-RF on-site looking at 1L14-4. Possible 2.5W amp replacement. [1305] 2.5 W amp changed. Experienced Hpa not ready faults, waiting for KHVO.

Downtime Incident CLOSED: MZA8R04

Log Number: 3819013Logbook(s): Body: IDDescriptionPeriodCause8268Title:
MZA8R04

Summary:
Mismatching and wouldn't come up after a cycle. Shunt module was reset.

Downtime Incident CREATED: MZA8R04

Log Number: 3819012Logbook(s): Body: IDDescriptionPeriodCause8268Title:
MZA8R04

Summary:
Mismatching and wouldn't come up after a cycle. Shunt module was reset.

CREW CHIEF SIGN IN

Log Number: 3819006Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 08-01-20
Time : 15:50:25
Day : Saturday
Shift : Swing

Crew Chief : Aiken
Program Deputy: Benesch
Team : F
Arm : Deir
Designated SSO: Forman
LERF Operator:
Operators : Forman, Deir

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

HYST AREA 6-2-1 Cycles Magnets

Log Number: 3819004Logbook(s): Body: 

HYST AREA Version 6-2-1 has successfully cycled the following:

Area.Arcs.Arc2.Spreader
Area.Arcs.Arc2.Arc
Area.Arcs.Arc2.Recombiner
Area.Arcs.Arc2.Extraction
Area.Arcs.Arc4.Spreader
Area.Arcs.Arc4.Arc
Area.Arcs.Arc4.Recombiner
Area.Arcs.Arc4.Extraction
Area.Arcs.Arc6.Spreader
Area.Arcs.Arc6.Arc
Area.Arcs.Arc6.Recombiner
Area.Arcs.Arc6.Extraction
Area.Arcs.Arc8.Spreader
Area.Arcs.Arc8.Arc
Area.Arcs.Arc8.Recombiner
Area.Arcs.Arc8.Extraction
Area.Arcs.ArcA.Spreader
Area.Arcs.ArcA.Arc

Box Supplies Were Cycled

Log Number: 3819002Logbook(s): Tags: Body: 

The following box supplies were CYCLED by version 2-7 of the Box Supply Cycling Script:
ARC2A Ramp Rate = 6.762 A/s
ARC4A Ramp Rate = 11.048 A/s
ARC6A Ramp Rate = 9.130 A/s
ARC8A Ramp Rate = 11.948 A/s
ARCAA Ramp Rate = 12.239 A/s
RSEP8A Ramp Rate = 10.471 A/s

Phaser Results

Log Number: 3818997Logbook(s): Body: Results where Job # is 354 {1} CAVITY PHASE ERROR (DEGREES) OUTCOME PHASE (DEGREES) START DATE DURATION (SECONDS) CORRECTION DATE CORRECTION ERROR REASON 1L14-4 10.73 CORRECTED -171.52 01-Aug-2020 15:21 57<

1L14-4 No HV after 2.5w AMP Replacement

Log Number: 3818991Logbook(s): Body: 

S. Holben and L. Farrish replaced 2.5w AMP at ~ 1330 and then the zone would not go back into HV. I came in to help Sam figure out why and found no ready enable on the CPS. It tried once after turning the AC Control Power and AC current three phase but dropped out. Tried the AC Control switch again to get a ready and the CPS made it into HV and stayed. 90 percent humidity and 85 degree heat affecting contacts in the Linacs again. Techs: C. Kerns, S. Holben and L. Farrish.

LEM North Linac

Log Number: 3818988Logbook(s): Tags: Body: LEM 9.6 was successfully run for the North Linac on Sat Aug 1 15:07:57 2020. The application phase took 0:00:12 to complete.

Pages